Session: Nov 2004 (session 3)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (FIocchi Franco at Thu Nov 4 08:39:27 2004) Success :
Trere were fog

Onsala (Michael Lindqvist at Thu Nov 4 09:26:39 2004) Success :
No known problems.

Effelsberg (A. Kraus at Thu Nov 4 10:35:32 2004) Success :

Jodrell_Bank (Alastair Gunn at Mon Nov 8 11:35:01 2004) Failure :
MkII was used. Mk2 could not observe due to ongoing engineering. No alternative 22GHz antenna available.

Cambridge (Alastair Gunn at Mon Nov 8 11:41:10 2004) Success :
No known problems.

Metsahovi (Ari Mujunen at Tue Nov 9 14:12:26 2004) Success :
No known problems.

Robledo (ZP for Esther Moll at Fri Nov 12 13:06:29 2004) Minor failures :
1) No sources were lost 2) Diode control is possible in K-band and full_cal procedure worked and noise diodes responded during the whole observation. You may find the tsys1 information available in the observation log. 3) During first scan a estrange behavior of the XPLOT window showing LCP data was observed. To be sure that it was only a monitoring problem the diodes were turned on and off and the LCP K-band receiver (MMS#1) was reconfigured. 4) Further analysis of the log showed that the LCP K-band receiver and all video converters worked ok. 5) The formatter remained in sync 6) Tape#2 recording started at 04:25:45z instead of 04:24:00z. There was not enough time for tape change. As a result, footage at the end of scan# 1120 was 16186 feet instead of 17600 feet.

Noto (P.Cassaro at Tue Nov 16 15:05:53 2004) Minor failures : WIND
From 17:06 to 19:06 UT the antenna has been put in stow due to high wind.